### v1

# Comments on “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques” (arXiv:2507.07180 [hep-ph])

Abstract
We give our comments on Ref. [1] which critiques our idea of exploring quantum statistics to distinguish between Dirac and Majorana neutrinos proposed in some of our earlier works. We show that the ad-hoc symmetrization of the Dirac case amplitude square advocated in Eqs. (16) and (35) of [1], have no physical basis and it leads to violation of lepton number in the the standard model. Therefore, we consider this symmetrization ‘by hand’ as incorrect and misleading. Moreover, we show that it is even unnecessary for the proof of pDMCT in context of final states containing a neutrino and antineutrino pair.
I What is the primary issue and where does it appear in [1]?
In Ref. [1] the authors aim to provide a critique of our papers [2, 3, 4] where we have explored the possibility whether quantum statistics could help in distinguishing Dirac and Majorana neutrinos. We note that in Ref. [1] the authors have tried to force agreement with the practical Dirac Majorana confusion theorem (pDMCT), while in our works [2, 3, 4] we have explored the domain of applicability of pDMCT in addition to providing a general proof of it in context of final states having neutrino antineutrino pair. We have used the standard quantum field theory in our papers [2, 3, 4] and our approach very well accommodates the previously existing literature.
The underlying basis of the main claim of Ref. [1] where they disagree with our papers, is contained within their Eqs. (33)-(38) of Sec. IV. While discussing the case of Dirac neutrinos, the authors of [1] contend that for the unobserved final states of neutrino and antineutrino one should symmetrize the amplitude squared with respect to the 4-momenta of the neutrino and antineutrino. Just below Eq. (34) of [1] they state that “To make a fair comparison with the Majorana case, as the neutrinos are not observed, it is imperative that both momentum assignments be the anti-neutrino are summed to construct the Dirac matrix-element squared”. This notion of symmetrizing at the probability level for Dirac neutrinos affects the inferences made in Ref. [1]. As explained below, we find such symmetrization “by hand” to be incorrect and misleading. Our comments below exclusively discuss all aspects of this particular issue found in Ref. [1]. It should be mentioned that although we primarily focus on their Sec. IV where they have critiqued our work on the decay [2], the above mentioned issue can also be found in Sec. III of [1] around Eqs. (14)-(16). Finally, we also point out the misinterpretation of the helicity amplitudes given in Eq. (28) of [3] by the authors of [1] in their Eqs. (39)-(43). We clarify these aspects as well.
II Our comments
Our comments are as follows.
1.
There is no fundamental principle or law of physics which requires that the amplitude square for a process having two distinguishable particles in the final state ought to be symmetrized with respect to the exchange of the 4-momenta of the concerned particles, when the two particles are not detected in the detector111We note that with advancement of technology, particle detection and identification also improves, and even with the best of detector technology some particles might not get detected by the detector either due to limited spatial coverage of the detector or due to the finite time needed for triggering events in a detector.. A Dirac neutrino is distinguishable from the corresponding antineutrino. Therefore, for a final state containing a pair of Dirac neutrino and antineutrino there is no need to do any symmetrization with respect to their momentum exchange.
Detection and identification of a particle in a detector does not affect the amplitude square which is a purely theoretical computation where one assumes that all the 4-momenta of all the particles are well known. The quantity which gets affected by experimental detection or observation or measurement is the “observable” (such as some asymmetry, or distribution, or differential decay rate, or scattering cross-section etc.) which involves both the amplitude square as well as the phase space considerations. Non-observation of any final particle is well taken care of by suitably performing corresponding phase space integration. We will later show how such phase space integration for invisible neutrinos, without any ‘by hand’ symmetrization of the Dirac amplitude square, directly leads to pDMCT which we have already provided in [2]. But before that we point out how such symmetrization is incorrect and misleading.
2.
In the standard model (SM) the Dirac neutrino () and antineutrino () carry lepton numbers as they are produced via different weak charged currents (CCs) and for , to ensure lepton number conservation within the SM. When considering Dirac neutrino antineutrino pair production via two CCs, as it happens in the decays considered in Ref. [1] as well as in our work [2], the neutrino and antineutrino 4-momenta are paired with the corresponding charged lepton so that lepton number is conserved. This is not required for Majorana neutrinos, in which and the lepton number is not a good (conserved) quantum number. However, adding a term to the Dirac case amplitude square where the neutrino and antineutrino 4-momenta have been exchanged, as done in the right sides of Eqs. (16) and (35) of Ref. [1], amounts to considering and both of which violate lepton number. When working with Dirac neutrinos in the framework of the SM alone without consideration of any new physics contributions, as is the case in decay considered in both [1] and [2], one should avoid considering lepton number violation effects.
3.
The authors of Ref. [1] state in the paragraph containing Eq. (32) that we have not been explicit about how we construct the Dirac as well as Majorana amplitudes in [2]. We note that we have explicitly and fully mentioned construction of our Dirac case and Majorana case amplitudes in Sec. IV, sub-sections A, B, C and D of [2]. As far as the spin or helicity of the neutrinos are concerned, they are automatically taken care of by the nature of the CC vertices. To elaborate slightly more on this let us first focus on the Dirac case, the 4-momenta of antineutrino and neutrino in [2] are denoted by and respectively. The amplitude in our case is denoted by and given by Eq. (15a) in [2].222The momentum-exchanged amplitude , is not at all required in the Dirac case. There is no underlying symmetry principle that connects the two amplitudes, and , in the general case. We then follow the standard procedure in any QFT calculation to evaluate the amplitude square in such a case, i.e. we take the modulus square of the amplitude, sum over final spins, average over initial spins333Since is a pseudo-scalar meson its spin is and no averaging over initial spins is needed., and evaluate the traces that appear during the calculation. We clarify that our calculation in [2] does not explicitly consider specific helicities of the neutrino and antineutrino.
We again reiterate that the problem of detectability of particles is still not relevant at the amplitude stage. As mentioned before, it is relevant only when computing any observable for experimental study, in which case we need to integrate over all the unobservable phase space variables so that the final observable thus obtained involves only measurable quantities. We emphasize, there is no formal reasoning for symmetrizing the amplitude squared as done in Ref. [1]. Even when required by quantum statistics, any symmetrization (for identical bosons) or antisymmetrization (for identical fermions) is done at the level of amplitude only. The question of detectibility can neither alter the rules for writing the transition amplitude nor the computation of the amplitude square.
4.
The authors of Ref. [1] wrongly presume that our calculation in [2] follows steps analogous to what they suggest in their Eqs. (40) and (43). As clarified above, the calculation in [2] does not explicitly consider any specific helicity states for the neutrino and antineutrino pair in the final state. The suggestion made in Eq. (28) of Ref. [3] is a generic suggestion where the decay amplitude can be written as a sum of a set of mutually orthogonal helicity amplitudes. Although we have not used the helicity amplitude formalism, as mentioned before and as we have noted below Eq. (30c) of [2], we have summed over all final spins in our calculations instead.
III Explanations regarding non-symmetrization of Dirac amplitude square and pDMCT
Here we explain how proving pDMCT does not require the ad-hoc symmetrization of Dirac case amplitude square as done in [1]. If there is no way to infer the energies or momenta of the neutrino and antineutrino experimentally, then any experimental observable, say for Dirac (D) or Majorana (M) neutrinos, would always involve integration over the full 4-momenta of neutrino and antineutrino. In such a case it does not matter whether we assign and to antineutrino and neutrino respectively or vice versa, because
Here act as dummy variables for integration as they both have same integration ranges. In the Majorana case, for the observable we get,
which implies that the observational difference between Dirac and Majorana case observables in such a case is proportional to ,
because the interference term here is always (except in the case of collinear neutrinos, i.e. for ) proportional to as it involves helicity flips. This is the most general way to prove pDMCT in context of processes that have a neutrino and antineutrino pair in their final states. Most importantly,
- 1.
2.
it does not require any by hand symmetrization of the Dirac case amplitude square as done in [1].
One of the most direct violations of pDMCT occurs when one considers the collinear neutrino antineutrino pair, (say). In such a case, considering the decay , it can be shown that Majorana amplitude vanishes in this case due to anti-symmetrization, the amplitude for the Dirac case is non-zero, i.e.
and as given in [2] for Dirac case we get
In Dirac case the probability is not only non-zero, but is also not proportional to , which defies the logic used in Ref. [1]. In fact it remains the same and is non-zero even if one wrongly symmetrizes the amplitude squared as done in [1].
Since quantum statistics of identical particles does not depend on the mass of the particle but its spin, one expects that there would be certain signatures of quantum statistics that explicitly shows this behavior, i.e. there would be some predictions from quantum statistical consideration which do not depend on mass. Our exploration of quantum statistics to distinguish between Dirac and Majorana neutrinos is based on this positive expectation and it is bolstered by the collinear case mentioned above although experimentally we do not yet have a way to study the collinear case. Moreover, it would be unwise to interpolate results of any particular process to a different process which might involve different interactions altogether. For example, in our study of (see Eq. (1) of [2], Eq. (16) of [3], Eq. (21) of [5]) we find that in the SM we always get
which is in confirmation of pDMCT even at the amplitude square level. In comparison, for doubly weak charged current mediated processes such as , when one does not use any ad-hoc symmetrization of Dirac amplitude square, we get from Eqs. (31), (32) of [2],
Then the challenge here is to find out an experimentally measurable quantity, a suitable observable, without completely destroying the finite difference between Dirac and Majorana cases. Our study has helped us realize that there are three major ways one might circumvent pDMCT,
1.
if direct and exchange amplitudes in Majorana case have non-trivially different contributions (Eq. (6) of [2] which gets satisfied in Eqs.(31) and (32)),
- 2.
- 3.
IV Conclusion
We have considered the claims made in Ref. [1] in context of our papers [2, 3, 4] and presented clarifications for considering them to be both invalid and incorrect. The main crux of the arguments presented in our papers [2, 3, 4] is to show that quantum statistics can not only distinguish between Majorana or Dirac types, but also define the domain of validity of pDMCT in the most general way possible in the absence of a formal proof of it until now. As everyone agrees, pDMCT is valid when we integrate over the momenta of neutrino and antineutrino (using the Dirac nomenclature). The proposed ad-hoc symmetrisation in [1] at the probability level for Dirac case simply achieves this by a different procedure, but as explained before we do not consider this approach to be correct. Nevertheless, our main focus in [2, 3, 4, 5, 6] is on the possible exceptions to pDMCT when we can fix the neutrino-antineutrino momenta either directly (possible in principle but difficult experimentally) or indirectly (as pointed out in [2] under special kinematic conditions) or when new physics effects are taken into consideration.
Acknowledgements.
The work of CSK is supported by grants from the National Research Foundation (NRF) of the Korean government (RS-2022-NR070836).References
- [1] I. Bigaran, S. J. Parke and P. Pasquini, “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques,” [arXiv:2507.07180 [hep-ph]].
- [2] C. S. Kim, M. V. N. Murthy and D. Sahoo, “Inferring the nature of active neutrinos: Dirac or Majorana?,” Phys. Rev. D 105, no.11, 113006 (2022) doi:10.1103/PhysRevD.105.113006 [arXiv:2106.11785 [hep-ph]].
- [3] C. S. Kim, “Practical Dirac Majorana confusion theorem: issues and applicability,” Eur. Phys. J. C 83, no.10, 972 (2023) doi:10.1140/epjc/s10052-023-12156-9 [arXiv:2307.05654 [hep-ph]].
- [4] C. S. Kim, D. Sahoo and K. N. Vishnudath, “Searching for signatures of new physics in to distinguish between Dirac and Majorana neutrinos,” Eur. Phys. J. C 84, no.9, 882 (2024) doi:10.1140/epjc/s10052-024-13262-y [arXiv:2405.17341 [hep-ph]].
- [5] C. S. Kim, J. Rosiek and D. Sahoo, “Probing the non-standard neutrino interactions using quantum statistics,” Eur. Phys. J. C 83, no.3, 221 (2023) doi:10.1140/epjc/s10052-023-11355-8 [arXiv:2209.10110 [hep-ph]].
- [6] G. Hernández-Tomé, C. S. Kim and G. L. Castro, Eur. Phys. J. C 85, no.6, 686 (2025) doi:10.1140/epjc/s10052-025-14379-4 [arXiv:2411.09124 [hep-ph]].