### v2

# Comments on “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques” (arXiv:2507.07180 [hep-ph])

Abstract
We give our comments on Ref. [1] which critiques the idea of using quantum statistics to distinguish between Dirac and Majorana neutrinos proposed in some of our earlier works [2, 3, 4]. The ad hoc symmetrization of the Dirac case amplitude square advocated in Eqs. (16) and (35) of [1] has no physical basis and it leads to violation of lepton number in the standard model for Dirac neutrinos. Therefore, this symmetrization ‘by hand’ is in principle incorrect.
I Note Added after publication of [1]
In the published version of the paper [1] we find essentially the same mistakes which we had pointed out before (see Sec. II for a concise version of our original comments).
1.
Their Dirac case amplitude square for the doubly weak charged decay of the meson: , as given in their Eq. (35) has two terms. The second term has been added by hand to their Eq. (34). The first and second terms of their Eq. (35) are related to each other via exchange of the neutrino antineutrino 4-momenta, . This makes their Dirac case amplitude square symmetric under , even though Dirac neutrino and antineutrino are distinguishable particles. The addition of the exchanged second term without simultaneously interchanging the muon 4-momenta, ie. , implies that the lepton number is violated for the second term of their Eq. (35). The anti-symmetrization of amplitude for Majorana neutrino is enforced by Pauli principle due to the identical nature of Majorana neutrino and antineutrino, and it is unrelated to detection or non-detection of the final neutrinos. Therefore, the ad hoc symmetrization of Dirac case amplitude square under exchange in Eq. (35) of [1] is incorrect.
2.
The Eqs. (16) and (35) of [1] seem to suggest that if one does not observe some of the non-identical final state particles, then one needs to symmetrize the amplitude square of the process by adding terms that exchange the unobserved 4-momenta. We find this as a rather strange claim. It is a well known fact that in any collider experiment some particles escape detection or identification, when they are either very close to the beam pipe or when the tracks get too crowded or if a part of the detector malfunctions or if many long-lived particles are present in the final state which decay outside the detector, etc. In such cases the amplitude square of the process is never symmetrized for non-detection of any of the distinct particles. Instead, the amplitude and amplitude square of the process are calculated assuming that all the non-identical particles are detectable and distinguishable in principle, and the non-detection of any particle is taken care of by appropriately constraining the phase space integral.
3.
In our paper [2], we have used a simplified notation to denote the amplitude of the doubly weak charged decay of the meson: . Our simplified notation for the Dirac case amplitude, as given in Eq. (2) of our paper [2] has the following expanded form,
where we have now explicitly mentioned all the 4-momenta and spins of the final particles. For brevity of expressions we chose to highlight only the dependence, while other dependencies were implicitly always there. As mentioned below our Eq. (3) in [2], the amplitude squares, e.g. , include summation over all final spins and averaging of initial spins. Note that since the parent particle has spin , the average over initial spins is trivial/meaningless in our case.
4.
Just above Eq. (43) of [1], the authors claimed that “… exactly vanishes when … .” This sentence is absolutely misleading. As explained above, the spin information is implicit in our simplified amplitude notation. Therefore, the correct statement would have been “… exactly vanishes when and … .”
II Summary of our Original Comments
In Ref. [1] the authors aim to provide a critique of our papers [2, 3, 4] where we have explored the possibility whether quantum statistics could help distinguish Dirac and Majorana neutrinos. We note that in Ref. [1] the authors have tried to force agreement with the practical Dirac Majorana confusion theorem (pDMCT), while in our works [2, 3, 4] we have explored the domain of applicability of pDMCT in addition to providing a general proof of it in context of final states having neutrino antineutrino pair. We have used the standard quantum field theory in our papers [2, 3, 4] and our approach very well accommodates the previously existing literature. Our comments are as follows.
1.
The underlying basis of the main claim of Ref. [1] where they disagree with our papers, is contained within their Eqs. (33)-(38) of Sec. IV, as well as in Eqs. (14)-(16) of their Sec. III. While discussing the case of Dirac neutrinos, the authors of [1] contend that for the unobserved final states of neutrino and antineutrino one should symmetrize the amplitude squared with respect to the 4-momenta of the neutrino and antineutrino. Just below Eq. (34) of [1] they state that “To make a fair comparison with the Majorana case, as the neutrinos are not observed, it is imperative that both momentum assignments be the anti-neutrino are summed to construct the Dirac matrix-element squared”. This notion of symmetrizing at the probability level for Dirac neutrinos affects the inferences made in Ref. [1].
There is no fundamental principle or law of physics which requires that the amplitude square for a process having two distinguishable particles in the final state ought to be symmetrized with respect to the exchange of the 4-momenta of the concerned particles, when the two particles are not detected in the detector111We note that with advancement of technology, particle detection and identification also improves, and even with the best of detector technology some particles might not get detected by the detector either due to limited spatial coverage of the detector or due to the finite time needed for triggering events in a detector.. A Dirac neutrino is distinguishable from the corresponding antineutrino. Therefore, for a final state containing a pair of Dirac neutrino and antineutrino there is no need to do any symmetrization with respect to their momentum exchange.
Detection and identification of a particle in a detector does not affect the amplitude square which is a purely theoretical computation where one assumes that all the 4-momenta of all the particles are well known. The quantity which gets affected by experimental detection or observation or measurement is the “observable” (such as some asymmetry, or distribution, or differential decay rate, or scattering cross-section etc.) which involves both the amplitude square as well as the phase space considerations. Non-observation of any final particle is well taken care of by suitably performing corresponding phase space integration. We have already pointed out in the beginning of Sec. III of our paper [2] as well as in Sec.2.2 of [3] that such phase space integration for invisible neutrinos directly leads to pDMCT, and this does not require any ad hoc symmetrization of the Dirac case amplitude square as done in [1].
In the standard model (SM) the Dirac neutrino () and antineutrino () carry lepton numbers as they are produced via different weak charged currents (CCs) and for , to ensure lepton number conservation within the SM. When considering Dirac neutrino antineutrino pair production via two CCs, as it happens in the decays considered in Ref. [1] as well as in our work [2], the neutrino and antineutrino 4-momenta are paired with the corresponding charged lepton so that lepton number is conserved. This is not required for Majorana neutrinos, for which and the lepton number is not a good (conserved) quantum number. However, adding a term to the Dirac case amplitude square where the neutrino and antineutrino 4-momenta have been exchanged, as done in the right sides of Eqs. (16) and (35) of Ref. [1], amounts to considering and both of which violate lepton number. When working with Dirac neutrinos in the framework of the SM alone without consideration of any new physics contributions, as is the case in decay considered in both [1] and [2], one should avoid considering lepton number violation effects.
2.
The authors of Ref. [1] state in the paragraph containing Eq. (32) that we have not been explicit about how we construct the Dirac as well as Majorana amplitudes in [2]. We note that we have explicitly and fully mentioned construction of our Dirac case and Majorana case amplitudes in Sec. IV, sub-sections A, B, C and D of [2]. As far as the spin or helicity of the neutrinos are concerned, they are automatically taken care of by the nature of the CC vertices. To elaborate slightly more on this let us first focus on the Dirac case, the 4-momenta of antineutrino and neutrino in [2] are denoted by and respectively. The amplitude in our case is denoted by and given by Eq. (15a) in [2].222The momentum-exchanged amplitude , is not at all required in the Dirac case. There is no underlying symmetry principle that connects the two amplitudes, and , in the general case. We then follow the standard procedure in any QFT calculation to evaluate the amplitude square in such a case, i.e. we take the modulus square of the amplitude, sum over final spins, average over initial spins333Since is a pseudo-scalar meson its spin is and no averaging over initial spins is needed., and evaluate the traces that appear during the calculation. We clarify that our calculation in [2] does not explicitly consider specific helicities of the neutrino and antineutrino.
We again reiterate that the problem of detectability of particles is still not relevant at the amplitude stage. As mentioned before, it is relevant only when computing any observable for experimental study, in which case we need to integrate over all the unobservable phase space variables so that the final observable thus obtained involves only measurable quantities. We emphasize, there is no formal reasoning for symmetrizing the amplitude squared as done in Ref. [1]. Even when required by quantum statistics, any symmetrization (for identical bosons) or antisymmetrization (for identical fermions) is done at the level of amplitude only. The question of detectibility can neither alter the rules for writing the transition amplitude nor the computation of the amplitude square.
3.
The authors of Ref. [1] wrongly presume that our calculation in [2] follows steps analogous to what they suggest in their Eqs. (40) and (43). As clarified above, the calculation in [2] does not explicitly consider any specific helicity states for the neutrino and antineutrino pair in the final state. The suggestion made in Eq. (28) of Ref. [3] is a generic suggestion where the decay amplitude can be written as a sum of a set of mutually orthogonal helicity amplitudes. Although we have not used the helicity amplitude formalism, as mentioned before and as we have noted below Eq. (30c) of [2], we have summed over all final spins in our calculations instead.
We have considered the claims made in Ref. [1] in the context of our papers [2, 3, 4] and presented clarifications to consider them both invalid and incorrect. The main crux of the arguments presented in our papers [2, 3, 4] is to show that quantum statistics can not only distinguish between Majorana or Dirac types, but also define the domain of validity of pDMCT in the most general way possible in the absence of a formal proof of it until now. We note that pDMCT is strictly valid when we integrate over the momenta of neutrino and antineutrino (using the Dirac nomenclature). The proposed ad hoc symmetrisation in [1] at the probability level for the Dirac case simply achieves this by a different procedure, but as explained before, we do not consider this approach to be correct. Nevertheless, our main focus in [2, 3, 4, 5, 6] is on the possible exceptions to pDMCT when we can fix the neutrino-antineutrino momenta either directly (possible in principle but difficult experimentally) or indirectly (as pointed out in [2] under special kinematic conditions) or when new physics effects are taken into consideration.
Acknowledgements.
The work of CSK is supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2022-NR070836, RS-2026-25471898). We thank the authors of [1], especially Innes Bigaran and Stephen Parke, for engaging in discussions on our paper that brings clarity to many of the issues.References
- [1] I. Bigaran, S. J. Parke and P. Pasquini, “Exploring Quantum Statistics for Dirac and Majorana Neutrinos using Spinor-Helicity techniques,” Phys. Rev. D 112, no.5, 053008 (2025) [arXiv:2507.07180 [hep-ph]].
- [2] C. S. Kim, M. V. N. Murthy and D. Sahoo, “Inferring the nature of active neutrinos: Dirac or Majorana?,” Phys. Rev. D 105, no.11, 113006 (2022) doi:10.1103/PhysRevD.105.113006 [arXiv:2106.11785 [hep-ph]].
- [3] C. S. Kim, “Practical Dirac Majorana confusion theorem: issues and applicability,” Eur. Phys. J. C 83, no.10, 972 (2023) doi:10.1140/epjc/s10052-023-12156-9 [arXiv:2307.05654 [hep-ph]].
- [4] C. S. Kim, D. Sahoo and K. N. Vishnudath, “Searching for signatures of new physics in to distinguish between Dirac and Majorana neutrinos,” Eur. Phys. J. C 84, no.9, 882 (2024) doi:10.1140/epjc/s10052-024-13262-y [arXiv:2405.17341 [hep-ph]].
- [5] C. S. Kim, J. Rosiek and D. Sahoo, “Probing the non-standard neutrino interactions using quantum statistics,” Eur. Phys. J. C 83, no.3, 221 (2023) doi:10.1140/epjc/s10052-023-11355-8 [arXiv:2209.10110 [hep-ph]].
- [6] G. Hernández-Tomé, C. S. Kim and G. L. Castro, Eur. Phys. J. C 85, no.6, 686 (2025) doi:10.1140/epjc/s10052-025-14379-4 [arXiv:2411.09124 [hep-ph]].